what is the substitution method of x+y=-2 and 2x-y=-4
Answer by ptaylor(2198)

what is the substitution method of x+y=-2 and 2x-y=-4
x+y=-2---------------------eq1
2x-y=-4---------------------eq2
When we are solving a system of equations by substitution, the idea is to find the easiest approach. There is no rigid method to solving equations by substitution.
For your problem, here are the possibilities:
From eq1:
y=-2-x--------------substitute -2-x for y in eq2
or
x=-2-y---------------substitute -2-y for x in eq2
From eq2:
y=4+2x---------------substitute 4+2x for y in eq1
or
x=(-4-y)/2-------------substitute (-4-y)/2 for x in eq1
Any of the above substitutions will yield the correct answer but it appears that either of the first two may be the easiest.
I'll do the last one. It appears to be the most difficult:
(-4-y)/2+y=-2 multiply each term by 2
-4-y+2y=-4 subtract 4 from each side
-4+4-y+2y=-4+4 collect like terms
y=0--------------substitute y=0 into eq1
x+0=-2;
x=-2
CK
from eq1:
-2+0=-2
-2=-2
from eq2
2*-2-0=-4
-4=-4
Hope this helps---ptaylor